OpenStudy (anonymous):

Find the sum of the first 60 terms in the arithmetic series 9+14+19+...+304 I think the sum of the first 60 terms is 9390 I want to know if that answer is correct

OpenStudy (oaktree):

The answer is n/2(first+last) = 30(9+304) = 9390. So yes.